Plantation Homes Company is considering the acquisition of Condominiums, Inc. early in 2015. To assess the amount it might be willing to pay, Plantation Homes makes the following computations and assumptions.

A. Condominiums, Inc. has identifiable assets with a total fair value of $14,379,000 and liabilities of $8,680,000. The assets include office equipment with a fair value approximating book value, buildings with a fair value 27% higher than book value, and land with a fair value 75% higher than book value. The remaining lives of the assets are deemed to be approximately equal to those used by Condominiums, Inc.
B. Condominiums, Inc.’s pretax incomes for the years 2012 through 2014 were $1,204,000, $1,598,000, and $1,026,000, respectively. Plantation Homes believes that an average of these earnings represents a fair estimate of annual earnings for the indefinite future. However, it may need to consider adjustments to the following items included in pretax earnings:
Depreciation on buildings (each year) 905,000
Depreciation on equipment (each year) 48,200
Extraordinary loss (year 2014) 318,000
Sales commissions (each year) 229,000
C.

The normal rate of return on net assets for the industry is 16%.

(a)

Assume further that Plantation Homes feels that it must earn a 26% return on its investment and that goodwill is determined by capitalizing excess earnings. Based on these assumptions, calculate a reasonable offering price for Condominiums, Inc. Indicate how much of the price consists of goodwill. Ignore tax effects.
(b)
Answer #1
Normal earnings on the industry=Net assets*normal rate of return on net assets
Net assets=Fair value of assets-Fair value of liabilities=14379000-8680000=$ 5699000
Normal earnings in the industry=5699000*16%=$ 911840
Adjusted earnings:
office equipment with a fair value approximating book value. Hence, depreciation will continue to be same for future years.No adjustment required for depreciatio on equipmet
Sales commission will also be same for future years. Hence, no adjustment required.
buildings with a fair value 27% higher than book value.Hence, depreciatio on building will have an effect o future earnings.
2012 2013 2014
Pre-tax income 1204000 1598000 1026000
Add: Extra ordinary loss 318000
Less:Additional depreciation on building
(905000*27%) -244350 -244350 -244350
Adjusted earnings 959650 1353650 1099650
Excess earnings:
$
Average of adjusted earnings (959650+1353650+1099650)/3 1137650
Less: Normal earnings in the industry 911840
Excess earnings 225810
Goodwill= Excess earnigs/Desired return on investment=225810/26%=$ 868500
Reasonable offering price=Net assets+goodwill=911840+868500=$ 1780340